OpenStudy (anonymous):

a card printer charges $5.00 to set up a job and additional $4.00 per box of 100 cards printed.what is the greastest number of boxes you could have printed for under $100

OpenStudy (anonymous):

100-5 = 95 95/4 = 23.75 =23 23*100 = 2300 cards

OpenStudy (anonymous):

or 23 boxes

OpenStudy (anonymous):

23 boxes? Sorry, I am not sure at all. I did it through this method--> 4x+5=100 ; x=95/4 which is equal to 23.75. So, the maximum number of boxes is 23 and cards = 23*100 which is equal to 2300.